### v1

# Exponential optimization of quantum state preparation via adiabatic thermalization

Abstract
The preparation of a given quantum state on a quantum computing register is a typically demanding operation, requiring a number of elementary gates that scales exponentially with the size of the problem. Using the adiabatic theorem for state preparation, whose error decreases exponentially as a function of the thermalization time, we report an explicit analytic expression for the dependence of the characteristic time on the Hamiltonian used in the adiabatic evolution. Exploiting this knowledge, we then design a preconditioning term that modifies the adiabatic preparation, thus reducing its characteristic time and hence giving an exponential advantage in state preparation. We prove the efficiency of our method with extensive numerical experiments on prototypical spin-models, which gives a promising strategy to perform quantum simulations of manybody models via Trotter evolution on near-term quantum processors.
Introduction.– We are currently living the so-called Noisy Intermediate-Scale Quantum (NISQ) era [1], in which actual devices are made of a limited number of noisy qubits available in a number of different technological platforms. Restricting to the realm of digital quantum processors satisfying the DiVincenzo criteria [2], typically 30 all-to-all connected qubits are available in trapped ions quantum processors [3] and more than 100 in superconducting quantum circuits with next-neighbor coupling [4]. The physics of manybody quantum systems may thus be efficiently simulated through the direct application of such devices [5, 6]. A proven advantage in quantum simulations is yet to be presented, at time of writing. In fact, in principle one expects NISQ hardware to overcome known issues in simulating quantum systems on classical devices, such as the sign problem for fermionic systems [7, 8]. Nowadays, a substantial limitation resides in the severely restricted hardware resources; nevertheless first proof-of-concept quantum simulations of chemical properties have been reported [9, 10, 11], as well as the time-dynamics of manybody Hamiltonians [12, 13, 14] and quantum field theories [15, 16, 17, 18, 19].
In quantum simulations on NISQ devices, one of the most pressing targets is the efficient generation of a desired quantum state of a given model Hamiltonian on a quantum register, which is essential, e.g., to directly compute expectation values of the observables under study. This is the focus of the present work. We define $H_{1}$ as the target Hamiltonian, properly mapped on a quantum computer, and denote with $\{|\Omega_{n}(1)\rangle\}$ the set of its eigenstates. Using a sequence of single- and two-qubit gates [20], a unitary operator $U_{n}$ acting on a register of $L$ qubits such that $|\Omega_{n}(1)\rangle=U_{n}|0\rangle^{\otimes L}$ may always be found, and the number of gates required to exactly implement $U_{n}$ scales exponentially with $L$, i.e., with the size of the system. Currently, the noisy character of state-of-the art qubits and quantum gates available in NISQ devices limits the number of operations that can be reliably executed on actual hardware. Hence, significant effort has been devoted to propose efficient algorithms allowing to obtain a well approximated $U_{n}$ within limited quantum computing resources. For state preparation, variational quantum eigensolvers have been proposed as a valid solution, consisting of hybrid quantum/classical algorithms in which quantum circuits with parametrized operations are classically optimized to achieve the desired state [21, 22, 23, 24]. Despite their usefulness for simple models, a general strategy to define such circuits for arbitrary systems is still not available; moreover, they may involve demanding classical optimizations in high-dimensional spaces [25]. An alternative approach has been recently proposed [26], based on imaginary time evolution, which might prove more effective than variational eigensolvers on NISQ devices, despite its intrinsic limitations in the preparation of quantum states with long correlation lengths.
Here we take a different route, by applying the Adiabatic Thermalization (AT) [27] to show that any eigenstate of a target Hamiltonian can be generated, in principle, with high fidelity and limited quantum computing resources. The key idea is to start from a Hamiltonian $H_{0}$, different from $H_{1}$, whose eigenstates are analytically known and trivial to prepare on a quantum register. Then, its $n$-th eigenstate $|\Omega_{n}(0)\rangle$ is evolved in time from $t=0$ to $t=\tau$, according to the auxiliary Hamiltonian
$\small{H\left(s\right)=\left(1-f\left(s\right)\right)H_{0}+f\left(s\right)H_{1}}\,,\quad s=t/\tau\,,$ (1)
which is parametrized in terms of a dimensionless variable $s\in[0,1]$. Essentially, $H\left(s\right)$ interpolates between $H_{0}$ and $H_{1}$ through the time-dependent holomorphic function $f(s)$ such that $f(0)=0$ and $f(1)=1$. In what follows we denote the time evolution operator associated with $H(s)$ in the time interval $[0,\tau]$ as $U_{\tau}$. Its application to the n-th eigenstate of $H_{0}$ produces the final state $U_{\tau}\ket{\Omega_{n}(0)}$, which is an approximation of the actual eigenstate |$\ket{\Omega_{n}(1)}$ of $H_{1}$, with a discrepancy quantified by the infidelity
$\mathcal{I}(\tau)\equiv 1-\absolutevalue{\bra{\Omega_{n}(1)}U_{\tau}\ket{\Omega_{n}(0)}}^{2}\,.$ (2)
The present work is based on the observation that an appropriate choice of $H_{0}$, or equivalently $\Delta=H_{1}-H_{0}$, may lead to computational speed up. Starting from known analytical results on the exponential convergence of this procedure [29], we hereby focus on the analytic evaluation of the error in quantum state preparation as a function of the thermalization time, which we then use to formulate a novel strategy to select $H_{0}$ such that the residual thermalization effect is further reduced when computing the approximate quantum state. To support our formal result, we show the effectiveness of this approach by explicitly analysing the Heisenberg model in either one or two dimensions, and demonstrating by numerical simulations the suppression of residual thermalization effects at a fixed quantum circuit depth. Given the known limitations of the AT to prepare singly targeted eigenstates [28], we hereby extend our protocol also to the preparation of eigenspaces spanned by more than one eigenstate. These results are considerably relevant for prospective quantum simulations of manybody models on NISQ hardware with unprecedented system size, possibly towards quantum advantage.
Exponential scaling.– Let $\sigma_{0}(s)$ be a subset of the full spectrum $\sigma(s)$ of $H(s)$, and $\mathcal{H}$ the Hilbert space of the system configurations. We address the problem of preparing a state belonging to the subspace $V(s)\subset\mathcal{H}$ spanned by those eigenstates with eigenvalues in $\sigma_{0}(s)$. Let $P(s)$ be the projector on $V(s)$, our aim is then to efficiently evolve the subspace $P(0)\mathcal{H}$ into $P(1)\mathcal{H}$ through AT. In what follows, the thermalization error is quantified as the operator norm
$\epsilon_{\mathrm{AT}}(\tau)\equiv\norm{\left(\mathds{1}-P(1)\right)U_{\tau}P(0)}\,,$ (3)
which is the generalization of Eq. (2) to the case of $\mathrm{dim}\left(V(s)\right)>1$, where the norm of an operator $A$ acting on the Hilbert space $\mathcal{H}$ is defined as
[…]
Figure 1: Schematic illustration of the adiabatic evolution of the spectral features related to a time-dependent Hamiltonian $H(s)$, in which $\sigma(s)$ is the full spectrum, partitioned into the subsets of eigenvalues $\sigma_{0}(s)$ and $\sigma_{1}(s)$. Here, $D(s)$ denotes the bandwidth of $\sigma_{0}(s)$, while $d(s)=\mathrm{min}\left[d_{+}(s),d_{-}(s)\right]$ defines the minimal gap between the two subsets of eigenvalues.
For finite-dimensional Hilbert spaces, the parameter $D$ is inherently finite, which guarantees the exponential falloff of the error just by the conditions that $H(s)$ is holomorphic and $d(s)>0$. By defining the characteristic decay time as
$\tilde{g}\geq\mathrm{sup}_{s\in[0,1]}g(s)\,,$ (6)
the exponential bound
$\epsilon_{\mathrm{AT}}\leq C\,\mathrm{exp}\left(-\left\lfloor\tau/\tilde{g}\right\rfloor\right)$ (7)
is straightforwardly obtained. Clearly, any reduction of $\tilde{g}$ exponentially reduces the error bound in Eq. (7). The purpose of this work is then to find an analytic expression for $g(s)$ that will ultimately allow us to define an efficient optimization strategy. In particular, we hereby report our extension of the result of Ref. [29] which relates $g(s)$ to the metrics $\{\norm{\Delta},D(s),d(s)\}$, therefore providing a concrete way to control $\tilde{g}$. For the purposes of the present manuscript, it is relevant to notice that these metrics can be efficiently estimated by classical means, e.g., via the Lanczos Algorithm [31, 32]. Given the functions
$W(d,D)=\left[2^{8}\left(1+2\frac{D}{\pi d}\right)^{2}\left(\frac{\pi^{2}}{3}+\frac{4}{d}\right)\right]^{7/3}\,,$ (8)
and
$\varrho\left(s,\tfrac{d}{\norm{\Delta}}\right)=\mathrm{sup}_{\rho\in\mathbf{C}}\left[|\rho|:\left|f(s+\rho)-f(s)\right|<\tfrac{d}{4\norm{\Delta}}\right]$ (9)
in the case of $H(s)$ defined according to Eq. (1), we obtain that the ratio
$g(s)=4\frac{W\left(d(s),D(s)\right)}{\varrho\left(s,\frac{d(s)}{\norm{\Delta}}\right)}$ (10)
satisfies Eq. (5) (our proofs will be reported elsewhere). Since $g(s)$ is a non-increasing function of $d(s)$ and $1/D(s)$, we fix $\tilde{g}$ to
$\tilde{g}=4\frac{W(d_{\mathrm{min}},D_{\mathrm{max}})}{\mathrm{inf}_{s\in[0,1]}\varrho(s,\frac{d_{\mathrm{min}}}{\norm{\Delta}})}\,,$ (11)
where
$d_{\mathrm{min}}=\mathrm{inf}_{s\in[0,1]}d(s)\,,\quad D_{\mathrm{max}}=\mathrm{sup}_{s\in[0,1]}D(s)\,.$ (12)
For the commonly employed linear holomorphic function $f(s)=s$, the expression above reduces to
$\tilde{g}=16\frac{\norm{\Delta}}{d_{\mathrm{min}}}W(d_{\mathrm{min}},D_{\mathrm{max}}))\,.$ (13)
Given a target Hamoltonian $H_{1}$, the naive choice for $H_{0}$ is the diagonal part of $H_{1}$, such that its eigenstates are manifest. In what follows we present a universal method to select a more convenient auxiliary Hamiltonian, $H_{0}$, in the context of digital quantum simulations, where $U_{\tau}$ is constructed using the Trotter-Suzuki approximation [33]. Our proposal may not lead to the absolute best choice for specific problems, but it guarantees an exponential advantage, in the quantum state preparation of a target Hamoltonian $H_{1}$, over the naively picked $H_{0}$. The most generic starting Hamiltonian can be expressed in terms of its matrix elements with respect to the computational basis as
[…]
in which $M$ denotes a parameterized Hermitian operator (called the preconditioner) to be optimized in order to minimize $\tilde{g}$. We impose three conditions on $M$: (i) $H_{0}$ should be diagonal, such that its eigenstates are the elements of the computational basis and it is then trivial to initialize them on a quantum computing register; (ii) the additional computational cost for the implementation of $\mathcal{U}_{M}(\theta)=\mathrm{exp}\left(-i\theta M\right)$ at each Trotter step (where $\theta\in\mathbf{R}$) should not frustrate the cost reduction stemming from the minimization of $\tilde{g}$; finally, (iii) the number of real parameters needed to describe $M$ should not increase exponentially with the number of qubits, $L$. The first condition is satisfied by any diagonal operator. We hereby propose the extremely “cheap” form
$\bra{i}M\ket{i}=\Sigma_{j=0}^{L-1}\alpha_{j}\,\mathrm{bin}_{j}\left(i\right)\,,$ (15)
in which |$\ket{i}$ is the i-th element of the computational basis, and $\mathrm{bin}_{j}\left(i\right)$ is the j-th digit of $i$ once expressed in binary notation. The key idea behind the choice of Eq. (15) will be reported elsewhere. Most importantly, this choice allows to fulfill both conditions (ii) and (iii). In fact, the number of free parameters is $L$ in this case, thus increasing linearly with the system size. This result is essential to reduce the quantum computational cost of this procedure. In particular, the corresponding time evolution operator is
$\mathcal{U}_{M}(\theta)=\bigotimes_{j=0}^{L-1}\mathrm{P}_{j}(\alpha_{j}\theta)\,,$ (16)
in which $\mathrm{P}_{i}$ is the single-qubit phase gate, applied to the i-th qubit. Notice that the quantum circuit depth of $\mathcal{U}_{M}(\theta)$ corresponds to a single qubit gate, independently from the size of the system.
Refer to captionRefer to caption
Figure 2: Ground state preparation for the 9 sites 1D and 2D Heisenberg models with periodic boundary conditions, schematically illustrated in panels (a) and (b), respectively, where the error dependence is plotted as a function of thermalization time, $\tau$, for $J^{x}=5J^{z}$.
Essentially, we propose to interpolate between $H_{0}$ (Eqs. (14)-(15)) and the targeted $H_{1}$ by AT, after classically minimizing $\tilde{g}$ through a scanning of the free parameters $\{\alpha_{j}\}$ in $H_{0}$. This simple protocol will guarantee an exponentially faster quantum state preparation.
Figure 3: Characteristic times $g_{\mathrm{fit}}$ obtained from fitting the numerical curves $\epsilon_{\mathrm{AT}}(\tau)$ (as the ones in Fig. 2) with a single exponential (see text), for different values of the ratio $J^{x}/J^{z}$ of the 1D Heisenberg model. Standard deviations of the fit are plotted as error bars.
Numerical results.– Here we apply the AT to the preparation of the ground state for the Heisenberg model in either one- or two-dimensional (1D, 2D) square lattices with periodic boundary conditions. The purpose is to numerically confirm the previous theoretical result in case models of practical relevance for quantum simulations. The corresponding quantum circuits are implemented and executed on the Qiskit qasm simulator [34].
To focus only on $\epsilon_{\mathrm{AT}}$, we use sufficiently short Trotter steps, to suppress the systematic error coming from the Trotter-Suzuki approximation. In the following, the state preparation is performed by applying three different approaches, all with the same adiabatic function $f(s)=s$. The first is the standard AT, where $M$ is set to zero in Eq. (14). Since the calculation of $\norm{\Delta}$ is computationally cheap and independent from the chosen target subspace $V(s)$, we consider a second approach where the parameters $\alpha_{j}$ are chosen to minimize $\norm{\Delta}$ in Eq. (13). Notice that this does not guarantee that $\tilde{g}$ is minimized. In fact, we finally consider a third approach consisting in the full classical evaluation of $\tilde{g}$ in Eq. (13) involving the estimation $d_{\mathrm{min}}$ and $D_{\mathrm{max}}$. Despite yielding a significantly better scaling for our case studies, for more complicated models it may be less practical, since this approach requires the knowledge of the spectrum at several intermediate values of $s$.
The XZ Heisenberg model is a prototypical manybody model describing systems of spin-1/2 particles on a lattice, defined by the interacting Hamiltonian
$H_{1}=-\frac{1}{2}\Sigma_{\langle i,j\rangle}\left(J^{z}Z_{i}Z_{j}+J^{x}X_{i}X_{j}\right),$ (17)
where the sum runs over next-neighboring sites, while $X_{i}$ and $Z_{i}$ are the Pauli matrices acting on the i-th qubit, conventionally defined as
$X=\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix}\,,\quad Y=\begin{bmatrix}0&-i\\ i&0\\ \end{bmatrix}\,,\quad Z=\begin{bmatrix}1&0\\ 0&-1\\ \end{bmatrix}\,.$ (18)
In this work we consider real and positive coupling constants, $J^{z}$ and $J^{x}$. The AT process is then described by the operator
$H(s)=-\frac{1}{2}\Sigma_{\langle i,j\rangle}\left(J^{z}Z_{i}Z_{j}+sJ^{x}X_{i}X_{j}\right)+(1-s)M\,.$ (19)
At $s=0$ the Hamiltonian is diagonal and, in the unpreconditioned case ($M=0$), has a degenerate ground state. The corresponding eigenstates, |$\ket{0}^{\otimes L}$ and |$\ket{1}^{\otimes L}$, are orthonormal and invariant under translations. It is possible to preserve the translational invariance of the AT process by imposing the condition $\alpha_{i}=\alpha$ ($\forall$ $i$) in Eq. (16). This choice, albeit not unique, allows to reduce the number of parameters to a single one. The effect of the preconditioner ($M$) on the physical system is equivalent to the introduction of an external uniform magnetic field along the $z$ axis, with field intensity $\alpha$. The results for the ground state preparation of this model are reported in Fig. 2, for either the 1D or the 2D models. The outcome is quite striking: independently on the dimensionality of the model or the parameters regime, the AT provides a way of determining the ground state of the XZ Heisenberg model with an error that scales almost exponentially with the thermalization time. In addition, by comparing the three different approaches above, we notice a significant reduction in the required thermalization time to achieve a given error, with the third (most optimal) approach allowing to obtain the shortest $\tau$. The performance differences between the different approaches become more and more evident on increasing $J^{x}/J^{z}$, as shown in Fig. 3. Here, the characteristic time obtained by fitting “$\ln{\{\epsilon_{\mathrm{AT}}(\tau)\}}$ )” with the function “$C_{\mathrm{fit}}-\tau/g_{\mathrm{fit}}$” is reported as a function of $J^{x}/J^{z}$. In particular, the $g_{\mathrm{fit}}$ values obtained with the three approaches are nearly indistinguishable when $J^{x}\ll J^{z}$, while they exhibit a clear advantage after applying our preconditioning strategy for $J^{x}\gg J^{z}$, i.e. when the target Hamiltonian is significantly different from the starting one. The peak in the plot occurring around $J^{x}/J^{z}\simeq 1$ corresponds to the critical point where, for an infinite number of spins, the system would exhibit a phase transition. In that region the preparation of the ground state is still a rather expensive task and our proposal is less effective. Similar results have been obtained for the Ising model in a transverse magnetic field, in both one and two dimensions, and will be reported elsewhere.
Conclusions.– We have applied the theory of Adiabatic Thermalization (AT) to the framework of quantum simulations on NISQ devices, specifically focusing on the efficient preparation of a given quantum state. Starting from the seminal result of Nenciu [29], we have reported an exponential boundary for the AT error (analytical proof reported elsewhere), with an explicit expression for its characteristic time in terms of quantities that can be classically estimated. Using this improved analytic control, we propose a universal preconditioning term, denoted as $M$ and tailored for NISQ computing, which can be effectively optimized by using the aforementioned formulation. Upon incorporation into the Hamiltonian model to be simulated on quantum hardware, $M$ allows to exponentially suppress the AT error. Finally, we have tested our protocol by numerical simulations performed for the representative Heisenberg and Ising models (only the former were explicitly shown), in either one or two dimensions. Our results demonstrate an exponential improvement, particularly notable under more strenuous computational scenarios. Despite performing numerical tests on simple spin models, we expect similar behaviours to occur for more complicated systems. We expect that our work will significantly enhance the efficiency of quantum state preparation, thereby boosting the prospective achievement of quantum utility in state-of-art digital quantum devices.
Acknowledgments. This research was supported by the Italian Ministry of Research (MUR) through NRR project CN00000013 - National Research Center on “HPC, Big Data and Quantum Computing” and PNRR project PE0000023 - National Quantum Science Technology Institute (NQSTI). The research of M.B. for this work was partially funded through the MUR program for young researchers “Rita Levi Montalcini”. The authors warmly acknowledge G. Guarnieri, F. Scala, and F. Tacchino for several scientific discussions.